# AI Governance through Markets

### v1

@@ A.7. Adversarial Testing
Adversarial testing assesses the resilience of AI systems by exposing them to challenging inputs designed to exploit vulnerabilities. These methods are essential for evaluating how AI models withstand malicious attempts, especially in security-sensitive applications (Ruiz et al., 2022). They help identify risks like data poisoning, evasion attacks, and extraction threats, which can compromise model integrity (Hannon et al., 2024).
@@ Appendix B Existing Disclosure Frameworks
Lack of corporate transparency coupled with inadequate risk assessment by investors creates a precarious foundation for potential catastrophic outcomes. The 2008 collapse of Lehman Brothers serves as a reminder about the far-reaching consequences of inadequate information transparency and (public, investors’, and regulatory) oversight in the markets. Leading up to its bankruptcy, Lehman had amassed significant exposure to subprime mortgages, which were increasingly seen as high-risk investments. As homebuyers defaulted on loans they couldn’t afford, it sent shockwaves to Wall Street. This led to a global financial crisis and millions of job losses and home foreclosures. Despite warning signs, such as declining housing prices and rising mortgage delinquencies, many investors continued to back Lehman’s stock. This disaster could’ve been uncovered at a nascent stage with rigorous oversight and due diligence. Lehman also employed deceitful accounting practices known as ”Repo 105,” temporarily removing debt from its balance sheet - a clear information asymmetry and lack of transparency between Lehman’s management and stakeholders.

### v2

@@ Acknowledgements.
We extend our deepest gratitude to Kristian Ronn, Ilan Strauss, Jamie Bernardi, Colleen McKenzie, Gabor Szorad and Alex Krasdomski for their contributions and invaluable insights to this research. We also thank Miles Brundage, Miguel Diaz, Nicolas Moes, Ema Provic, James Fox, Cyrus Hodes, Milan Griffes, Deric Cheng, Jacob Goodwin, Gabriel Weil, Lucy Philippon, Risto Uuk, Ryan Kidd, Shakeel Hashim, Deger Turan, Edmund Zagorin, David Lawrence and Oliver Klingfjord for their helpful support and feedback throughout this work.
@@ A.7. Adversarial Testing
Adversarial testing assesses the resilience of AI systems by exposing them to challenging inputs designed to exploit vulnerabilities. These methods are essential for evaluating how AI models withstand malicious attempts, especially in security-sensitive applications (Ruiz et al., 2022). They help identify risks such as data poisoning, evasion attacks, and extraction threats, which can compromise model integrity (Hannon et al., 2024).
@@ Appendix B Existing Disclosure Frameworks
Lack of corporate transparency coupled with inadequate risk assessment by investors creates a precarious foundation for potential catastrophic outcomes. The 2008 collapse of Lehman Brothers serves as a reminder about the far-reaching consequences of inadequate information transparency and (public, investors’, and regulatory) oversight in the markets. Leading up to its bankruptcy, Lehman had amassed significant exposure to subprime mortgages, which were increasingly seen as high-risk investments. As homebuyers defaulted on loans they couldn’t afford, it sent shockwaves to Wall Street. This led to a global financial crisis and millions of job losses and home foreclosures. Despite warning signs, such as declining housing prices and rising mortgage delinquencies, many investors continued to back Lehman’s stock. This disaster could have been uncovered at a nascent stage with rigorous oversight and due diligence. Lehman also employed deceitful accounting practices known as ”Repo 105,” temporarily removing debt from its balance sheet - a clear information asymmetry and lack of transparency between Lehman’s management and stakeholders.